OpenStudy (anonymous):

Find an equation of the tangent line to the hyperbola ((x^2)/(a^2))-((y^2)/(b2))=1 at the point (x0, y0).

OpenStudy (anonymous):

(b^2)(x^2)-(a^2)(y^2)=(a^2)(b^2) find y' 2b^2x-2a^2yy'=0 y'=(b^2x)/(a^2y) y'=m use point slope form (y-y0)=m(x-x0)